HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, Page 5 of 7  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  54   Tue Apr 1 22:53:59 2025 EstherGeneralLassie exports   2025-04-01_22-53-17-548_tcl1146.png2025-04-01_22-53-21-680_tcl1146.tdf 
  53   Tue Apr 1 22:05:24 2025 Ruijiu Analysisanalysis resultThis is a preliminary result during 4:00-19:00
April 1st 2025. 
I correct the frequency drift of 194Pb82+
 202504012157.png202504012157.pdf202504012206.pngphoto_2025-04-01_22-43-41.jpg 
  52   Tue Apr 1 22:03:07 2025 Jeroen, Zac, Esther, Shahab, WolframDAQOptimisation of settingsEstimation of potential gain by changing.
Files from this points are going to rubbish.
Potential gains are c. 20% 
  
  51   Tue Apr 1 21:05:20 2025 Jeroen, Zac, Esther, Shahab, WolframDAQChange of settings for shorter acq. window> Acquisition stopped at 20:36. 

> First file from this setting: RSA01-2025.04.01.03.52.59.247.tiq
  
  50   Tue Apr 1 20:37:04 2025 JeroenDAQChange of settings for shorter acq. windowAcquisition stopped at 20:36. 

First file from this setting: RSA01-2025.04.01.03.52.59.247.tiq
  
  49   Tue Apr 1 20:19:23 2025 Jeroen, Zac, EstherAcceleratorBeam fell out> > No beam from 19:57:41.

> Beam back from 20:02
 2025-04-01_20-17-10-651_tcl1064.png 
  48   Tue Apr 1 20:02:19 2025 Jeroen, Zac, EstherAcceleratorBeam fell out> No beam from 19:57:41.

Beam back from 20:02
  
  47   Tue Apr 1 20:00:14 2025 Jeroen, Zac, EstherAcceleratorBeam fell outNo beam from 19:57:41.   
  46   Tue Apr 1 17:08:13 2025 JG, ID, MC, CGDetectorspleiades in positionThe pleiades detectors are now moved into
a position which does not disturb the stored
beam. 
  
  45   Tue Apr 1 16:47:32 2025 Ruijiu, Jan, ShahabAnalysisThe frequency of 194Pb82+ is changed to 408.600 MHz   Screenshot_2025-04-01_16-46-56.pngphoto_2025-04-01_16-54-20.jpg 
  44   Tue Apr 1 16:44:32 2025 Jeroen, EstherGeneralNo SIS beam since 14:35Beam resumed at 16:40.   2025-04-01_16-47-13-677_tcl1146.png2025-04-01_16-47-10-870_tcl1146.tdf 
  43   Tue Apr 1 15:38:50 2025 ruijiu chenAnalysisfrequency drift   Screenshot_2025-04-01_15-38-26_1.pngScreenshot_2025-04-01_15-41-47.png 
  42   Tue Apr 1 15:32:59 2025 ZacDAQNTCAP RestartedNTCAP Has been restarted (we restart every
day around 15:00 so data goes in a new folder)
  
  41   Tue Apr 1 14:54:24 2025 allGeneralNo SIS beam since 14:35    
  40   Tue Apr 1 14:25:21 2025 ruijiu AnalysisThe frequency of 194Pb82+ has been changed from 408.606MHz to 408.602MHy during last 12 hoursThe following images display the frequency
spectrum at the start of every hour from
3:00 to 14:00.
 12x 
  39   Tue Apr 1 14:23:01 2025 JGGeneralPleiades pocket position testsWe started playing with the detector pockets
of the pleiades arrays at 14:23 01.04.2025.
  
  38   Tue Apr 1 14:17:07 2025 RuijiuAnalysisPID of 194Pb82+ 2025.04.01.14.13.07datafile = "/data.local/G22-00018_00203/results/analyzers/RSA01/RSA01-2025.04.01.14.13.07.068.tiq_spectrum.npz"
filep = "/data.local/G22-00018_00203/scripts/OnlineDataAnalysisSystem/rionid_exercises/exercise_194Pb_Exp_2025/New_TEline2022-ESR-194Pb-208Pb-nostripper_524MeV.lpp"
alphap = "1.399"
 Screenshot_2025-04-01_14-16-42.png 
  37   Tue Apr 1 14:02:23 2025 Jelena B, SiddharthGeneralLassie   2025-04-01_14-01-37-464_tcl1064.png 
  36   Tue Apr 1 12:58:27 2025 Jelena B, SiddharthGeneralLassie   2025-04-01_12-57-50-275_tcl1064.png 
  35   Tue Apr 1 11:47:14 2025 Jelena B, SiddharthGeneralLassie 2025-04-01_11-46-35-517_tcl1064.png    2025-04-01_11-46-35-517_tcl1064.png 
ELOG V3.1.5-fc6679b